Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
menzione - 6:59 Monday 15 April 2019 (45637) Print this report
Operator Report - Night Shift

ITF found in Locking. It reached LN3_SQZ at 21:05 UTC. Science Mode set at 21:06 UTC.
Science Mode for the rest of the shift with Horizon ~49 Mpc except for few drops (plot).

Guard Tours
21:30 UTC - 22:03 UTC
00:06 UTC - 00:35 UTC
02:29 UTC - 02:59 UTC
04:15 UTC - 04:43 UTC

Images attached to this report
Comments to this report:
swinkels - 11:43 Monday 15 April 2019 (45642) Print this report
I compared some spectra during and after the step at 1:46 UTC. No changes can be seen in the spectra of the main LSC signals and in Hrec_hoft_raw, while Hrec_hoft does see a significant change, see fig 1. One other thing to note is that there are some big drops in the horizon at the beginning of every step, see fig 2. One possibility could be that the big glitches somehow interfere with the measurement of some calibration lines, which then causes one of the noise subtraction schemes to be wrong for a few minutes. I looked several Hrec_OG* channels, but I didn't see any change, but maybe I am missing the right channel. To be confirmed by calibration experts ...

If true, the steps might be avoided by understanding and resolving the source of the big glitches, or by implementing some protection against outliers in the procedure that decides on the gain of the noise subtraction scheme.
Images attached to this comment
swinkels - 17:22 Monday 15 April 2019 (45644) Print this report
According to Benoit, the transfer functions for the subtractions are updated when the gps time modulo 500 seconds is zero. The steps seen in the horizon seem indeed to follow this pattern, see attached figure.
Images attached to this comment
Search
×

Warning

×